OpenStudy (anonymous):

Area of a circle with a diameter of 28.I keep getting 87.

Directrix (directrix):

What value did you use for the radius?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

A=pir^2 where r=d/2 so r=14.05 thus A=197.4pi (rough rounding)

OpenStudy (anonymous):

WHAT the answers are 88,300,600 and 800

OpenStudy (anonymous):

lmao maybe bc your not using units, there is a dimensional error... is everything is one unit?

OpenStudy (anonymous):

i Ddont know i think so

OpenStudy (anonymous):

Um Im thinking 88 but i think im wrong

OpenStudy (anonymous):

?

Directrix (directrix):

>> circle with a diameter of 28.I keep getting 87. I think you may be doing circumference.

OpenStudy (anonymous):

how do i change that

Directrix (directrix):

I need to know what the problem is as it appears in your text. I don't know what you mean by changing that. The statement of the problem tells you what to find. We cannot change the problem.